#3. In the freshmen class, 62% of the students take Math, 49% take Science, and 38% take both Science and Math. What percentage take at least one Science or Math course?
Answer #1

Let A be the event that a student takes math, B be the event that a student takes science. Then we are given that

P(A) = 62% =0.62 , P(B) = 49% =0.49 and P(A\bigcap B ) = 0.38 . We have to calculate the percentage of students that take at least one science or math course. In probability theory, at least one is given by A\bigcupB. Hence we have to calculate P(A\bigcupB). Using probability formula of set theory P(A\bigcupB) is given by

P(A\bigcupB) = P(A) + P(B) - P(A\bigcap B ) = 0.62+0.49 -0.38 = 0.73

Hence 73% students take at least one science or math course.